OpenStudy (anonymous):

Find E(X) , Var(X) , SD(X) if a random variable X is given by its density function f(x), such that f(x) = 0, if x<=0 f(x) = 3x^2, if 0 < x <= 1 f(x) = 0, if x > 1

OpenStudy (amistre64):

variance, standard deviation, and what is "E(x)"?

OpenStudy (anonymous):

i think it is the math expectation

OpenStudy (amistre64):

do we have a density function given or is this general?

OpenStudy (anonymous):

I am not sure the question is worded the same as i received it

OpenStudy (anonymous):

the density function is f(x)

OpenStudy (amistre64):

f(x) is a generic term meaning "any given function" so its asking for a generalization

OpenStudy (anonymous):

yes. I will assume that it is general. I do not know where to start

OpenStudy (amistre64):

0, if x<=0 f(x) = 3x^2, if 0 < x <= 1 0, if x > 1 this is saying that the f(x) is defined for a given domain; if the random variable is less than or equal to 0; then the probability = 0. If the random variable is greater than 1; then the probability = 0; otherwise, it is defined as 3x^2

OpenStudy (anonymous):

ok i get that part

OpenStudy (amistre64):

i believe this means that our distribution curve is parabolic and defined between 0 and 1; but i havent come across this b4 to be certain

OpenStudy (anonymous):

would there be 2 point that I would need for E(x)?

OpenStudy (amistre64):

i cant really say, i still dont have a clear understanding of what E(x) stands for :)

OpenStudy (amistre64):

expected value ..... ok

OpenStudy (amistre64):

expected value = mean of x*P(x)

OpenStudy (amistre64):

we got any choices to guide is with?

OpenStudy (amistre64):

do we know if our random variables are discrete or continuous?

OpenStudy (anonymous):

according to my notes it is the mathematical expectation of a continuous random variable

OpenStudy (amistre64):

if i read this right: E(x) = \(\int_{0}^1\ 3x^2\ dx\)

OpenStudy (amistre64):

well, I forgot an "x" in that :)

OpenStudy (amistre64):

3x^3 integrates to (3/4)x^4; at x=1 E(x) = 3/4 if my assumption is right

OpenStudy (amistre64):

and since E(x) is also the mean, we can see about determining the variance and sd with it

OpenStudy (amistre64):

the variance is equal to: \[\sum_{x=0}^{1}\ [(x-\bar x)^2P(x)]\]

OpenStudy (amistre64):

(0-.75)^2 * 3(0)^2 + (1-.75)^2 *3(1)^2 = .1875 ; but got no clue if thats right ....

OpenStudy (amistre64):

and of course the sd is just the square root of variance

OpenStudy (anonymous):

some kind of way we are supposed to have an additional E(X^2) equation

OpenStudy (anonymous):

then I think subtact them. so would the answer to E(x^2) be 3/5?

OpenStudy (amistre64):

that resembles the "short cut" sd in my book: \[\sqrt{\frac{n(\sum(x^2))-(\sum(x))^2}{n(n-1)}}\]

OpenStudy (amistre64):

perhaps: \[var(x)=\sum[(x^2)*3x^2] - E(x)\]??

OpenStudy (amistre64):

E(x)^2 that is

OpenStudy (anonymous):

ok now im confused lol

OpenStudy (anonymous):

i think i understand the main part of the equation though

OpenStudy (amistre64):

for continuous v random variable that wouold amount to: \(\int3x^4dx\) from 0 to 1 which equals \(\frac{3}{5}x^5\) from 0 to 1 which does = 3/5

OpenStudy (amistre64):

3/5 - (3/4)^2 = variance ?

OpenStudy (amistre64):

\[\frac{3}{5}-\frac{9}{16}=\frac{48-45}{80} = \frac{3}{80}\]

OpenStudy (anonymous):

I think thats correct it seems right. maybe i did it right. thank you

OpenStudy (amistre64):

:) good luck with it all